OpenStudy (anonymous):

please can someone help me write the equation of the circle in standard form x^2+y^2+4x+4y+4=0

OpenStudy (luigi0210):

Do you know how to complete the square?

OpenStudy (luigi0210):

First group the terms together and move constants to the other side \[(x^2+4x )+(y^2+4y )=-4\]

OpenStudy (luigi0210):

Can you handle it from here or so I need to continue?

OpenStudy (anonymous):

(x+2)^2 + (y^2+2)2=4 correct?

OpenStudy (luigi0210):

\[(x+2)^2+(y+2)^2 =4\] If this is what you mean then yes that is correct

OpenStudy (anonymous):

Thank you very much!

OpenStudy (luigi0210):

No Problem